OpenStudy (anonymous):

HELP<3 Medalllllllls Find the equation for the parabola with vertex (0, 0) and directrix x = 4.

OpenStudy (anonymous):

*breathes* Rather than fiddling with standard forms it's best to go back to basic focus-directrix definition. i.e. “dist from focus" = "perp dist from directrix” Let P(x,y) be any point on parabola then dist from focus (0,0) is √(x²+y²) perp dist from directrix x+y−4=0 is |x+y−4| / √(1²+1²) = |x+y−4| / √2 (this is a standard result, |...| is necessary because dist may come out ±) ∴ parabola is √(x²+y²) = |x+y−4| / √2 The √ and |...| can be eliminated by squaring both sides x²+y² = (x+y−4)² / 2 2(x²+y²) = x²+y²+16+2xy−8x−8y x²−2xy+y²+8x+8y−16 = 0

OpenStudy (anonymous):

Wow, thanks chica but the 4 answer choices are A. x = (1/16)y2 B. y = (1/16)x2 C. x = (–1/16)y2 D. y = (–1/16)x2

OpenStudy (anonymous):

You are amazingly educated though. You're gonna get far in life :3

OpenStudy (anonymous):

B. y = (1/16)x2 I'm Asian. LOL

OpenStudy (anonymous):

You're gorgeous, smart, and not american. What everyone wishes they were like.

OpenStudy (anonymous):

Lol. Thanks ^^